OpenStudy (anonymous):

There were 10 more two steppers than 5 times the number of waltzers. also there were 10 times as many two steppers as waltzers. how many of each?

OpenStudy (espex):

Something does not add up. If there are 10 more x than 5 times y would give you that x=10+5y Next it says that there are 10 times x as y giving you y=10x So when you substitute you get x=10+5(10x) which when worked out gives you that x=-10/49 and y=-100/49 This does not make sense.

OpenStudy (anonymous):

Solve the following for ts, the number of "two steppers" and w, the number of "waltzers".\[\left\{\text{ts}\text{ = }5 w +10,\frac{\text{ts}}{w}=10\right\}\]\[\{\text{ ts}\to 20,w\to 2\} \]